How fast is the surface area of a spherical balloon increasing when the radius is 10 cm and the volume is increasing at 15 cm3 /sec?


1
Expert's answer
2021-05-07T14:14:02-0400

Solution:

"\dfrac{dV}{dt}=15 cm^3/sec,\ r=10 cm, \dfrac{dA}{dt}=?"

"V=\dfrac43\pi r^3"

On differentiating w.r.t "r",

"\dfrac{dV}{dt}=4\pi r^2 \dfrac{dr}{dt}
\\ \Rightarrow 15=4\pi (10)^2 \dfrac{dr}{dt}
\\ \Rightarrow \dfrac{dr}{dt}=\dfrac{3}{80\pi} cm/sec"

Now, "A=4\pi r^2"

On differentiating w.r.t "r",

"\dfrac{dA}{dt}=8\pi r \dfrac{dr}{dt}=8\pi(10)(\dfrac{3}{80\pi})=3 cm^2/sec"